OpenStudy (anonymous):

differentiate 7e^x^2

OpenStudy (myname):

Here you just use the genral derivative to find e^x For example \[\int\limits_{}^{}e ^{x}dx = e ^{x}(x)\prime = e ^{x}(1)\] So in this case \[\int\limits_{}^{}7e ^{x ^{2}}dx=7\int\limits_{}^{}7e ^{x ^{2}}dx \] 7 is a constant Now, \[7\int\limits_{}^{}e ^{x ^{2}}dx=7e ^{x ^{2}}(x ^{2})\prime\] \[7e ^{x ^{2}}(x ^{2})\prime = 7e ^{x ^{2}}2x\] Final answer, \[\int\limits_{}^{}7e ^{x ^{2}}dx = 14xe ^{x ^{2}}\]

OpenStudy (anonymous):

14X*e^[x^2] final ans this is how u do it ....

OpenStudy (anonymous):

7e^(x^2)(x^2)' that prime mean deff ==== 14*2x*e^[x^2] u remember [x^2]'=2x? that is good... that is how u do it

OpenStudy (anonymous):

the first ans myname did is intergration but your question ask for differentiation

OpenStudy (myname):

I am very sorry that i mistakenly added the integration to it. It is like jmes said. Thanks jmes

OpenStudy (myname):

\[(7e ^{x ^{2}})\prime=7e ^{x ^{2}}(x ^{2})\prime\] \[7e ^{x ^{2}}(x ^{2})\prime=7e ^{x ^{2}}(2x)=14x e ^{x ^{2}}\]